If 10 grams of butane(C4H10), afuel in cigarrete lighter, react with 10 grams of O2, how many grams of the product will be formed

1
Expert's answer
2022-04-26T12:16:02-0400

2C4H10 + 13O2 → 8CO2 + 10H2O, n(O2)=0.3125 mol: n(C4H10)=0.1724 mol, the C4H10 is in excess, we have to calculate the mass of products wit using the n of O2.

m(products) = 18*((0.3125*10)/13) + 44*((0.3125*8)/10)=15.327 g of products will be formed